# Generalized Eckardt Points on del Pezzo surfaces of degree 1

### v1

@@ 3. Configurations of ten concurrent lines
A naive first step to give an upper bound for the answer to Question 1.1 is to count all cliques of size 10 in the graph $G$ on the 240 exceptional curves on a del Pezzo surface of degree 1. Although this does give an upper bound for the number of generalized Eckardt points on a del Pezzo surface of degree 1, it is very large (only counting those cliques with edges of weights 1 and 2 we obtain already 107,700,096 cliques of size 10, which is the sum of the orbits in Table 1), and far from sharp. For example, we know that some cliques in $G$ never correspond to exceptional curves that are concurrent in a point. A first step in answering Questions 1.1 and 1.2 is therefore to determine which cliques to count, that is, we want to determine in which way 10 concurrent lines on a del Pezzo surface of degree 1 can be concurrent. From Lemma 2.9 we know that exceptional curves can intersect each other with multiplicities 1, 2, or 3. Moreover, from Remark 2.10 we know how to distinguish between lines that are concurrent in a point on the ramification curve of the map (1), and lines that are concurrent in a point outside the ramification curve. In Theorem 3.3 we list all cliques that possibly correspond to concurrent exceptional curves. Moreover, we show that all cliques of the same isomorphism type form one orbit under the action of the Weyl group on the set of these cliques.
@@
We say that a clique $K$ in $G$ can be realized as a generalized Eckardt point, or simply realized, if there exists a del Pezzo surface of degree 1 with an isomorphism between the intersection graph on its 240 exceptional curves and $G$, such that $K$ corresponds to 10 concurrent exceptional curves under this isomorphism.
@@
  • (ii) Outside characteristic 2, there is only one orbit of cliques of size 10 in $G$ under the action of the Weyl group with only edges of weights 1 and 3 that can be realized as a generalized Eckardt point. Cliques in this orbit are of the form $\{c_{1},d_{1},\ldots,c_{5},d_{5}\}$, where $c_{i}\cdot c_{j}=1$ for $i\neq j$, and $c_{i}\cdot d_{i}=3$ for all $i$.
@@
(i). Recall that cliques in $G$ with only edges of weights 1 and 2 correspond to cliques in $\Gamma_{\{-1,0\}}$. The table in [WvL21, Appendix A] gives the complete list of orbits of the maximal cliques in $\Gamma_{\{-1,0\}}$, and we see that there are 6 orbits of maximal cliques of size 10. Translating this to cliques in $G$, a direct computation by hand shows that these are orbits $1-3$, $5-7$ in Figure 2. We are left with determining the orbits of the non-maximal cliques of size 10 in $G$ with only edges of weights 1 and 2. By [vLW23, Proposition 19], there are no maximal cliques in $\Gamma_{\{-1,0\}}$ of size $11$, and there is one orbit of maximal cliques of size $12$ under the action of $W_{8}$, which is the maximal size of cliques in $\Gamma_{\{-1,0\}}$. When translating this to the graph $G$, a maximal clique of size 12 in this orbit has the form
@@
Now that we have proved that there are 8 orbits of cliques of size 10 in $G$ with only edges of weights 1 and 2, we are left to show that in characteristic 0, at most 6 of them can be realized as a generalized Eckardt point. In [vLW23, Proposition 4.6] it is shown that a clique of type 4 can not be realized as a generalized Eckardt point outside characteristic 3, and the same is shown for a subset of a clique of type 7 in [DW, Proposition 5.4] in characteristic 0. From Lemma 3.2 we conclude that no cliques of type 4 or 7 can be realized as a generalized Eckardt point in these characteristics. This concludes part (i).
@@
Figure 2 contains the isomorphism types of 8 cliques in $G$ with only edges of weights 1 and 2. Cliques are fully connected subgraphs of $G$ with edges drawn corresponding to edges of weight 2. The edges not pictured in Figure 2 correspond to edges of weight 1 in the clique.
@@
Figure 2. Isomorphism types of 8 orbits.
@@
Example 3.6.
In [SvL14, Example 4.1], the authors construct a family of del Pezzo surfaces of degree 1 that all contain 10 concurrent exceptional curves, whose intersection graph is isomorphic to clique 5 in Figure 2. In [vLW23, Example 5.2] the authors construct a del Pezzo surface of degree 1 over a field of characteristic unequal to 2 with 10 exceptional curves that are concurrent on the ramification curve. We extend this to a whole family of such examples in Section 5.
Apart from clique 5 mentioned in Example 3.6, to our knowledge there is no example in the literature of 10 concurrent exceptional curves on a del Pezzo surface of degree 1 whose intersection graph is isomorphic to one of cliques 1, 2, 3, 6, or 8 in Figure 2. In Section 4 we discuss strategies for finding such examples, or for proving that more cliques are not realizable.
We finish this section with a brief discussion on the relation between Question 1.1 and the maximal number of generalized Eckardt points on del Pezzo surfaces of degree 2, a connection that was suggested to us by Bianca Viray. Over an algebraically closed field, every del Pezzo surface of degree 1 is the blow-up of a (non-unique) del Pezzo surface of degree 2, so another step in determining an upper bound for the number of generalized Eckardt points on a del Pezzo surface of degree 1 is to use what we know about del Pezzo surfaces of degree 2. A first question to answer is, which of the cliques in Figure 2 correspond to exceptional curves on a del Pezzo surface of degree 1 that, after blowing down an exceptional curve, could form a generalized Eckardt point on a del Pezzo surface of degree 2? In the following proposition we answer this question.
@@
Assume that the set $\{e_{1},\ldots,e_{10}\}$ contains a subset of 4 curves $e_{i_{1}},e_{i_{2}},e_{i_{3}},e_{i_{4}}$ that correspond to 4 concurrent exceptional curves on $Y$ after blowing down $c$. Since the exceptional curve on $Y$ are exactly the images of the 56 exceptional curves on $X$ that are disjoint from $c$, it follows that $e_{i_{1}},e_{i_{2}},e_{i_{3}},e_{i_{4}}$ are all disjoint from $c$. Choose a basis $[L],[E_{1}],\ldots,[E_{8}]$ for the exceptional classes in Pic $X$ as in Section 2. We can write every exceptional curve as a vector $(a,b_{1},\ldots,b_{8})$ with respect to this basis as denoted in the discussion below Theorem 2.4. Using this notation, observe that the set $S$ of 10 exceptional curves
[…]
$(4,2,1,2,1,2,1,1,1),(4,1,2,2,1,1,1,1,2),(4,1,1,2,1,1,2,2,1),(4,1,2,1,2,1,2,1,1),$
(4,2,1,1,2,1,1,2,1),(4,2,1,1,1,1,2,1,2)}(4,2,1,1,2,1,1,2,1),(4,2,1,1,1,1,2,1,2)\}( 4 , 2 , 1 , 1 , 2 , 1 , 1 , 2 , 1 ) , ( 4 , 2 , 1 , 1 , 1 , 1 , 2 , 1 , 2 ) }
@@
From Proposition 3.7 it follows that concurrent exceptional curves on a del Pezzo surface of degree 1 whose intersection graph looks like graphs 4, 5, or 8 in Figure 2 do not come from generalized Eckardt points on a del Pezzo surface of degree 2 (by blowing up a point). Therefore, we can not deduce an upper bound for the number of generalized Eckardt points realized by such cliques by using upper bounds for generalized Eckardt points on del Pezzo surfaces of degree 2. From Theorem 3.3 we know that clique 4 in Figure 2 can never be realized. For clique 8 this is not known, but we do know that clique 5 can be realized; see Example 3.6. It could therefore still be the case that a generalized Eckardt point on a del Pezzo surface of degree 1 can never come from a generalized Eckardt point of degree 2; for example, if it turns out that no clique except clique 5 in Figure 2 can be realized as a generalized Eckardt point.
@@ 4.1. Clique Representatives
Since $\ell_{7,8}$ contains $P$, but is distinct from the other three lines, then the line determined by $\ell_{7,8}$ must be of the form $x=gy$ for some $g\in k\setminus\{0\}$. Thus we can fix the following nine points for point set-up A
$P_{1}$ =(0:1:1)=(0:1:1),= ( 0 : 1 : 1 ) , $P_{5}$ =(1:1:1)=(1:1:1),= ( 1 : 1 : 1 ) ,
@@
An example of a del Pezzo surface of degree 1 with a generalized Eckardt point is an example of eight points in general position that pass through a given 10 concurrent exceptional curves. In order to determine whether such an example exists, we must determine whether it is possible to find eight points in general position of the above form so that these points determine 10 exceptional curves that pass through $P$. Since these eight points lie on lines $\ell_{1,2},\ell_{3,4},\ell_{5,6}$, and $\ell_{7,8}$ by construction, which are concurrent in $P$, we only need to check 6 more exceptional curves.
@@ 4.2. Strategy for Computing Examples
A general equation of a conic in $\mathbb{P}^{2}$ is given by $t_{1}x^{2}+t_{2}xy+t_{3}xz+t_{4}y^{2}+t_{5}yz+t_{6}z^{2}.$ Observe that a conic $b_{i,j,k}$ is determined by five points of $P_{1},P_{2},\ldots,P_{8}$ and that we want this conic to pass through $P$. Plugging the points in $\{P_{1},\ldots,P_{8},P\}\setminus\{P_{i},P_{j},P_{k}\}$ into the general equation for the conic gives a system of six linear homogeneous equations in six unknowns. At least one $t_{i}\neq 0$, so without loss of generality assume that $t_{6}\neq 0$. We solve the first equation for $t_{1}$ and substitute it in to the others. Then we solve the second equation for $t_{2}$ and substitute, continuing this process until we have a single equation of the form $Ft_{6}=0$ where $F\in k(a,b,c,d,e,f,g,h)$. Since $t_{6}\neq 0$, then $F=0$.
@@
Since $t_{3}\neq 0$, in order for all six points to lie on this conic, we need $de-d^{2}-e+1=0$. Notice that in order to solve for $t_{5}$ in the last step, we divided by $c-1$. This solution could be incorrect if $c=1$, but in this case $c=1$ is impossible since the points $P_{1},\ldots,P_{8}$ should be in general position.
@@ 5. Surfaces with generalized Eckardt points on the ramification curve
Let $k$ be any field of characteristic 00. For $a,b,c,d,e,f\in k$, define the following eight points in $\mathbb{P}_{k}^{2}$ as in point set-up A.
$P_{1}$ =(0:1:1)=(0:1:1),= ( 0 : 1 : 1 ) , $P_{5}$ =(1:1:1)=(1:1:1),= ( 1 : 1 : 1 ) ,
@@
We disregard the factor $s$, since we are interested in configurations of $P_{1},\ldots,P_{8}$ in general position such that $c_{2,1}$ contains $P$. Write
@@
We claim that, if $P_{1},\ldots,P_{8}$ are in general position, we have $p(b,c,d,e,f)\neq 0$ and we can write
$a=-q(b,c,d,e,f)/p(b,c,d,e,f).$
@@
Writing $a=-q(b,c,d,e,f)/p(b,c,d,e,f)$, we compute with magma the polynomial equation that corresponds to $c_{8,7}$ containing $P$, and we find that it factors as $F_{2}\cdot t=0$, where $t$ is a product of elements in $S$, and
@@
The intersection of $V_{1}$ with the hyperplanes $b=-1$, $c=\tfrac{5}{4}$, and $d=-1$ gives an elliptic curve of rank 1 in $\mathbb{A}^{5}$, that contains infinitely many points corresponding to a del Pezzo surface of degree 1 with a generalized Eckardt point.
@@ 6. Finite Fields
Theorem 6.1.
[vLW23][Theorem 1.1] Let $P\in X(k)$ be a point on the ramification curve of $\varphi$ in (1). The number of exceptional curves that go through $P$ is at most ten if char $k\neq 2$, and at most sixteen if char $k=2$.
Theorem 6.2.
[vLW23][Theorem 1.2] Let $P\in X(k)$ be a point outside of the ramification curve of $\varphi$ in (1). The number of exceptional curves that go through $P$ is at most ten if char $k\neq 3$, and at most twelve if char $k=3$.
Further, van Luijk and Winter produced examples of del Pezzo surfaces of degree 1 to show that the upper bounds in these theorems are sharp except possibly in char $k=5$ outside the ramification curve. This leaves the following open question.
@@
Does there exist a del Pezzo surface of degree 1 in char $k=5$ with 10 concurrent lines outside the ramification curve?

### v2

@@ 3. Configurations of ten concurrent lines
A naive approach to give an upper bound for the answer to Question 1.1 is to count all cliques of size 10 in the graph $G$ on the 240 exceptional curves on a del Pezzo surface of degree 1. Although this does give an upper bound for the number of generalized Eckardt points on a del Pezzo surface of degree 1, it is very large (only counting those cliques with edges of weights 1 and 2 we obtain already 107,700,096 cliques of size 10, which is the sum of the orbits in Table 1), and far from sharp. For example, we know that some cliques in $G$ never correspond to exceptional curves that are concurrent in a point [vLW23, Propositions 3.6, 4.6]. A better first step in answering Questions 1.1 and 1.2 is therefore to determine which cliques to count, that is, we want to determine in which way 10 concurrent lines on a del Pezzo surface of degree 1 can be concurrent. From Lemma 2.9 we know that exceptional curves can intersect each other with multiplicities 1, 2, or 3. Moreover, from Remark 2.10 we know how to distinguish between lines that are concurrent in a point on the ramification curve of the map (1), and lines that are concurrent in a point outside the ramification curve. In Theorem 3.3 we list all cliques of size 10 that possibly correspond to concurrent exceptional curves. Moreover, we show that cliques of the same isomorphism type form one orbit under the action of the Weyl group on the set of these cliques.
@@
We say that a clique $K$ in $G$ can be realized as a generalized Eckardt point, or simply realized, if there exists a del Pezzo surface of degree 1 with an isomorphism between the intersection graph on its 240 exceptional curves and $G$, such that $K$ corresponds to a set of exceptional curves concurrent in a generalized Eckardt point under this isomorphism.
@@
  • (ii) Outside characteristic 2, there is only one orbit of cliques of size 10 in $G$ under the action of the Weyl group with only edges of weights 1 and 3 that can be realized as a generalized Eckardt point. Cliques in this orbit are of the form depicted in Figure 3: a set of vertices $\{c_{1},d_{1},\ldots,c_{5},d_{5}\}$, where $c_{i}\cdot c_{j}=1$ for $i\neq j$, and $c_{i}\cdot d_{i}=3$ for all $i$.
@@
(i). Recall that cliques in $G$ with only edges of weights 1 and 2 correspond to cliques in $\Gamma_{\{-1,0\}}$. The table in [WvL21, Appendix A] gives the complete list of orbits of the maximal cliques in $\Gamma_{\{-1,0\}}$, and we see that there are 6 orbits of maximal cliques of size 10. Translating this to cliques in $G$, a direct computation by hand shows that these are orbits 131-3, 575-7 in Figure 2. We are left with determining the orbits of the non-maximal cliques of size 10 in $G$ with only edges of weights 1 and 2. By [vLW23, Proposition 19], there are no maximal cliques in $\Gamma_{\{-1,0\}}$ of size $11$, and there is one orbit of maximal cliques of size $12$ under the action of $W_{8}$, which is the maximal size of cliques in $\Gamma_{\{-1,0\}}$. When translating this to the graph $G$, a maximal clique of size 12 in this orbit has the form
@@
Now that we have proved that there are 8 orbits of cliques of size 10 in $G$ with only edges of weights 1 and 2, we are left to show that in characteristic 0, at most 6 of them can be realized as a generalized Eckardt point. In [vLW23, Proposition 4.6] it is shown that a clique of type 4 can not be realized as a generalized Eckardt point outside characteristic 3, and the same is shown for a subset of a clique of type 7 in [DW, Proposition 5.4] in characteristic 0. From Lemma 3.2 we conclude that no cliques of type 4 or 7 can be realized as a generalized Eckardt point outside characteristic 3 and in characteristic 0, respectively. This concludes part (i).
@@
Figure 2. The isomorphism types of 8 cliques in $G$ with only edges of weights 1 and 2. Each clique is a fully connected subgraphs of $G$ with the edges drawn corresponding to edges of weight 2. The edges not drawn correspond to edges of weight 1 in the clique.
Figure 3. The isomorphism type of a clique of size 10 in $G$ with only edges of weights 1 and 3. The clique is a fully connected subgraph of $G$ with the edges drawn corresponding to edges of weight 3. The edges not drawn correspond to edges of weight 1 in the clique.
@@
To our knowledge there is no example in the literature of 10 concurrent exceptional curves on a del Pezzo surface of degree 1 whose intersection graph is isomorphic to one of cliques 1, 2, 3, 6, or 8 in Figure 2. In Section 4 we discuss strategies for finding such examples, or for proving that more cliques are not realizable.
We finish this section with a brief discussion on the relation between Question 1.1 and the maximal number of generalized Eckardt points on del Pezzo surfaces of degree 2, a connection that was suggested to us by Bianca Viray. Over an algebraically closed field, every del Pezzo surface of degree 1 is the blow-up of a (non-unique) del Pezzo surface of degree 2, so another step in determining an upper bound for the number of generalized Eckardt points on a del Pezzo surface of degree 1 is to use what we know about del Pezzo surfaces of degree 2. A first question to answer is, which of the cliques in Figure 2 correspond to exceptional curves on a del Pezzo surface of degree 1 that, after blowing down an exceptional curve, could form a generalized Eckardt point on a del Pezzo surface of degree 2? In the following proposition we give a partial answer to this question.
@@
Assume that the set $\{e_{1},\ldots,e_{10}\}$ contains a subset of 4 curves $e_{i_{1}},e_{i_{2}},e_{i_{3}},e_{i_{4}}$ that correspond to 4 concurrent exceptional curves on $Y$ after blowing down $c$. Since the exceptional curve on $Y$ are exactly the images of the 56 exceptional curves on $X$ that are disjoint from $c$, it follows that $e_{i_{1}},e_{i_{2}},e_{i_{3}},e_{i_{4}}$ are all disjoint from $c$. Choose a basis $L],[E_{1}],\ldots,[E_{8}]$ for the exceptional classes in Pic $X$ as in Section 2. We can write every exceptional curve as a vector $(a,b_{1},\ldots,b_{8})$ with respect to this basis as denoted in the discussion below Theorem 2.4. Using this notation, observe that the set $S$ of 10 exceptional curves
$\{(1,0,0,0,0,0,0,1,1),(1,0,0,1,1,0,0,0,0),(1,1,1,0,0,0,0,0,0),(1,0,0,0,0,1,1,0,0),$
(4,2,1,2,1,2,1,1,1),(4,1,2,2,1,1,1,1,2),(4,1,1,2,1,1,2,2,1),(4,1,2,1,2,1,2,1,1),(4,2,1,2,1,2,1,1,1),(4,1,2,2,1,1,1,1,2),(4,1,1,2,1,1,2,2,1),(4,1,2,1,2,1,2,1,1),
(4,2,1,1,2,1,1,2,1),(4,2,1,1,1,1,2,1,2)}(4,2,1,1,2,1,1,2,1),(4,2,1,1,1,1,2,1,2)\}
@@
From Proposition 3.6 it follows that concurrent exceptional curves on a del Pezzo surface of degree 1 whose intersection graph looks like graphs 4, 5, or 8 in Figure 2 do not come from generalized Eckardt points on a del Pezzo surface of degree 2 (by blowing up a point). Therefore, we can not deduce an upper bound for the number of generalized Eckardt points realized by such cliques by using upper bounds for generalized Eckardt points on del Pezzo surfaces of degree 2. From Theorem 3.3 we know that clique 4 in Figure 2 can never be realized. We will show that clique 8 can be realized over $\mathbb{F}_{19}$ (Theorem 6.7). We also know that clique 5 can be realized, as mentioned in the introduction. It could be the case that, outside characteristic 3, a generalized Eckardt point on a del Pezzo surface of degree 1 outside the ramification curve can never come from a generalized Eckardt point on a del Pezzo surface of degree 2; for example, if it turns out that no clique except cliques 5 and 8 in Figure 2 can be realized as a generalized Eckardt point.
@@ 4.1. Clique Representatives
Since $\ell_{7,8}$ contains $P$ and is distinct from the other three lines, then the line determined by $\ell_{7,8}$ must be of the form $x=gy$ for some $g\in k\setminus\{0\}$. Thus we can fix the following nine points for point set-up A
$P_{1}$ $=(0:1:1),$ $P_{5}$ $=(1:1:1),$
@@
An example of a del Pezzo surface of degree 1 with a generalized Eckardt point can be constructed by finding eight points in general position that pass through a given set of 10 concurrent exceptional curves and then blowing up these points. In order to determine whether such an example exists, we must determine whether it is possible to find eight points in general position of the above form so that these points determine 10 exceptional curves that pass through $P$. Since these eight points lie on lines $\ell_{1,2},\ell_{3,4},\ell_{5,6}$, and $\ell_{7,8}$ by construction, which are concurrent in $P$, we only need to check 6 more exceptional curves.
@@ 4.2. Strategy for Computing Examples
A general equation of a conic in $\mathbb{P}^{2}$ is given by $t_{1}x^{2}+t_{2}xy+t_{3}xz+t_{4}y^{2}+t_{5}yz+t_{6}z^{2}.$ Observe that a conic $b_{i,j,k}$ is determined by five points of $P_{1},P_{2},\ldots,P_{8}$. As we want this conic to pass through $P$, we plug the points in $\{P_{1},\ldots,P_{8},P\}\setminus\{P_{i},P_{j},P_{k}\}$ into the general equation for the conic and obtain a system of six linear homogeneous equations in six unknowns. There is at least one $i$ with $t_{i}\neq 0$; by fixing such an $i$, we can rewrite the system of equations to obtain a single equation of the form $Ft_{i}=0$ where $F\in k(a,b,c,d,e,f,g,h)$. Since $t_{i}\neq 0$, then $F=0$.
@@
Since $t_{3}\neq 0$, in order for all six points to lie on this conic, we need $de-d^{2}-e+1=0$. Notice that in order to solve for $t_{5}$ in the last step, we divided by c1c-1. This solution could be incorrect if $c=1$, but in this case $c=1$ is impossible since the points $P_{1},\ldots,P_{8}$ should be in general position.
@@ 5. Surfaces with generalized Eckardt points on the ramification curve
Let $k$ be any field of characteristic 0. For $a,b,c,d,e,f\in k$, define the following eight points in $\mathbb{P}_{k}^{2}$ as in point set-up A.
$P_{1}$ $=(0:1:1),$ $P_{5}$ $=(1:1:1),$
@@
We disregard the factor $s$, since we are interested in configurations of $P_{1},\ldots,P_{8}$ in general position such that $c_{7,8}$ contains $P$. Write
@@
We claim that, if $P_{1},\ldots,P_{8}$ are in general position and $F_{1}(a,b,c,d,e,f)=0$, then we have $p(b,c,d,e,f)\neq 0$ and we can write
a=q(b,c,d,e,f)/p(b,c,d,e,f).a=-q(b,c,d,e,f)/p(b,c,d,e,f).
@@
Writing a=q(b,c,d,e,f)/p(b,c,d,e,f)a=-q(b,c,d,e,f)/p(b,c,d,e,f), we compute with magma the polynomial equation that corresponds to $c_{8,7}$ containing $P$, and we find that it factors as $F_{2}\cdot t=0$, where $t$ is a product of elements in $S$, and
@@
The intersection of $V_{1}$ with the hyperplanes b=1b=-1, $c=\tfrac{5}{4}$, and d=1d=-1 gives an elliptic curve of rank 1 in $\mathbb{A}^{5}$, that contains infinitely many points corresponding to a del Pezzo surface of degree 1 with a generalized Eckardt point.
@@ 6. Finite Fields
Theorem 6.1 ([vLW23][Theorem 1.1]).
Let $P\in X(k)$ be a point on the ramification curve of $\varphi$ in (1). The number of exceptional curves that go through $P$ is at most ten if char $k\neq 2$, and at most sixteen if char $k=2$.
Theorem 6.2 ([vLW23][Theorem 1.2]).
Let $P\in X(k)$ be a point outside of the ramification curve of $\varphi$ in (1). The number of exceptional curves that go through $P$ is at most ten if char $k\neq 3$, and at most twelve if char $k=3$.
Further, van Luijk and the fourth author produced examples of del Pezzo surfaces of degree 1 to show that the upper bounds in these theorems are sharp except possibly in char $k=5$ outside the ramification curve. This leaves the following open question.
@@
Does there exist a del Pezzo surface of degree 1 in characteristic 5 with 10 concurrent lines outside the ramification curve?